# Disentangling Representations through Multi-task Learning

### v2

@@ Appendix B Theoretical Derivations
  • $\hat{\mathbf{Y}}(t):=g(\mathbf{Z}(t))\in[0,1]^{N_{\textit{task}}}:$ Output vector of the multi-task classification model at time $t$, where each $\hat{Y}_{i}(t)$ is a Bernoulli random variable estimating the conditional probability $\Pr\{y_{i}(\mathbf{x}^{*})=1\}$ given the noisy observations (via latent variable $\mathbf{Z}(t)$ – see Equation 12).
@@ B.2 Trilateration via Multiple Decision Boundaries
: We derived an optimal estimator of $\mathbf{x}^{*}$ (denoted $\hat{\mathbf{X}}(t)$) based on noisy i.i.d. measurements $\mathbf{X}(1),\dots,\mathbf{X}(t)\sim\mathcal{N}(\mathbf{x}^{*},\sigma^{2}\mathbf{I}_{D})$ in Lemma B.1. In Lemma B.2 we derived the equation for Bernoulli variable $\hat{Y}(t)$ to estimate a single classification output $y(\mathbf{x}^{*})$ based on the same noisy measurements via $\hat{\mathbf{X}}(t)$. Finally, we showed in Lemma B.3 that the uncertainty in $\hat{Y}(t)$ and the time $t$ is sufficient to determine the projection distance between the decision boundary and $\mu(t)=\text{mean}(\mathbf{X}(1),\dots,\mathbf{X}(t))$ via Equation 15.
Let $\hat{\mathbf{Y}}(t)$ denote the vector of classification estimates $\hat{Y}(t)$ from Equation 15. We now have the tools to prove our final result via trilateration. Much like distance information from cell towers can be used to trilaterate [^3] Trilateration differs from triangulation, and it is more frequently used in practice. Triangulation is when one has angle information w.r.t. the cell towers. Usually, this is not available – so one trilaterates their position [57]. This more closely matches our setting, where we just have distances information w.r.t. the decision boundaries and must determine the position. one’s position, we will leverage Lemma B.3 and use distances from decision boundaries $\{(\mathbf{c}_{i},b_{i})\}_{i\in[N_{\textit{task}}]}$ to constrain the positions.
@@
If $\mathbf{C}$ is full-rank, then $\hat{\mathbf{Y}}(t)$, $t$, $\mathbf{b}$, and $\mathbf{C}$ are sufficient to reconstruct the exact value of $\mathbf{\mu}(t)$, the mean of $\mathbf{X}(1),\dots,\mathbf{X}(t)$, which is also the optimal estimator for $\mathbf{x}^{*}$.
@@
We may prove this claim by providing an algorithm to reconstruct $\mathbf{\mu}(t)=\text{mean}(\mathbf{X}(1),\dots,\mathbf{X}(t))$ from $\hat{\mathbf{Y}}(t),\mathbf{C}$, and $t$. Invoke Lemma B.3 to compute the signed projection distance between $\mathbf{\mu}(t)$ and each decision plane $(\mathbf{c}_{i},b_{i})$. Let $\mathbf{k}=[k_{1},\dots,k_{M}]^{\top}$ where each $k_{i}$ corresponds to decision boundary $\mathbf{c}_{i}$. Then the mean $\mathbf{\mu}(t)$ must satisfy
@@
Thus, for full rank $\mathbf{C}$, we will have a uniquely determined $\mu(t)$ value. ∎
@@
Any system that optimally estimates $\hat{\mathbf{Y}}$ based on noisy measurements $\{\mathbf{X}(1),\dots,\mathbf{X}(t)\}$ must implicitly encode a representation of optimal estimator $\mu(t)=\text{mean}(\mathbf{X}(1),\dots,\mathbf{X}(t))$ in its latent state $\mathbf{Z}(t)$ if decision boundary matrix $\mathbf{C}$ is full rank.
@@
This follows from the data processing inequality. We begin with the following Markov chain:
$\mathbf{x}^{*}\to\{\mathbf{X}(1),\dots,\mathbf{X}(t)\}\to\mathbf{Z}(t)\to\hat{\mathbf{Y}}(t)\to\mu(t)$ (17)
Put more simply, we have $\mathbf{x}^{*}\to\mathbf{Z}(t)\to\mu(t)$. Applying the data processing inequality, we obtain
$I\big{(}\mathbf{x}^{*};\mathbf{Z}(t)\big{)}\geq I\big{(}\mathbf{x}^{*};\mu(t)\big{)}$ (18)
where $I(\cdot;\cdot)$ denotes the mutual information between two variables.
Since $\hat{\mathbf{X}}(t)$ is the optimal estimator of $\mathbf{x}^{*}$ given measurements $\{\mathbf{X}(1),\dots,\mathbf{X}(t)\}$, $I\big{(}\mathbf{x}^{*};\mu(t)\big{)}=H(\mu(t))$. Therefore $I(\mathbf{x}^{*};\mathbf{Z}(t))\geq H(\mu(t))$, implying that $I(\mathbf{Z}(t);\mu(t))=H(\mu(t))$. In other words, $\mathbf{Z}(t)$ must contain all the information of $\mu(t)$, and $\mu(t)$ must be expressible as a deterministic function of $\mathbf{Z}(t)$. ∎
@@
$\tilde{\mu}=(\mathbf{C^{\top}C})^{-1}\mathbf{C}^{T}(\mathbf{k}+\mathbf{b})$ (19)
Conveniently, if the estimation errors in sub-optimal $\hat{\mathbf{Y}}$ have a mean of zero, additional decision boundaries in $\mathbf{C}$ (e.g., beyond the minimum $D$ linearly independent boundaries) result in improved estimation of $\mathbf{x}^{*}$ by the central limit theorem, thus generalizing our results to sub-optimal estimators.
@@ B.3 Optimal Representation Theorem (General Case)
If $\mathbf{C}\in{N_{\textit{task}}\times D}$ is a full-rank matrix representing the decision boundary normal vectors in $\mathbb{R}^{D}$ and $N_{\textit{task}}\geq D$, then any optimal estimator of $\mathbf{y}(\mathbf{x}^{*})$ must encode an optimal representation of the latent variable $\mathbf{x}^{*}$ in its latent state $\mathbf{Z}(t)$. Furthermore, the mutual information between $\mathbf{x}^{*}$ and $\mathbf{Z}(t)$ is at least as great as the entropy of the theoretically optimal estimator $\mu(t)$, ensuring that all the information of $\mu(t)$ is contained in $\mathbf{Z}(t)$. Consequently, $\mu(t)$ – the optimal estimate of $\mathbf{x}^{*}$ based on $f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))$ – can be written as a deterministic function (Equation 20) of latent state $\mathbf{Z}(t)$.
@@
$\mathbf{x}^{*}\to\{\mathbf{X}(1),\dots,\mathbf{X}(t)\}\to\mathbf{Z}(t)\to\hat{\mathbf{Y}}(t)\to\mu(t).$ (21)
@@
Applying the data processing inequality, we have:
$I\big{(}\mathbf{x}^{*};\mathbf{Z}(t)\big{)}\geq I\big{(}\mathbf{x}^{*};\mu(t)\big{)},$ (22)
where $I(\cdot;\cdot)$ denotes the mutual information between two variables. Since $\hat{\mathbf{X}}(t)$ is the optimal estimator of $\mathbf{x}^{*}$ given measurements $\{\mathbf{X}(1),\dots,\mathbf{X}(t)\}$, $I\big{(}\mathbf{x}^{*};\mu(t)\big{)}=H(\mu(t))$. Therefore, $I(\mathbf{x}^{*};\mathbf{Z}(t))\geq H(\mu(t))$, implying that $I(\mathbf{Z}(t);\mu(t))=H(\mu(t))$. In other words, $\mathbf{Z}(t)$ must contain all the information of $\mu(t)$, and $\mu(t)$ must be expressible as a deterministic function of $\mathbf{Z}(t)$.
@@
where $g(\mathbf{Z}(t))$ represents the transformation from the latent state to the classification probabilities. This completes the proof by showing that $\mu(t)$ is a deterministic function of $\mathbf{Z}(t)$. ∎
Motivated by the similarity between $\Phi(z)$ and sigmoid-like activation functions $g(z)$, we show that the two can be approximately canceled in Equation 20, implying that $\mu(t)$ can be reconstructed with high accuracy with a linear-affine transformation (e.g., linear decoding) when $g=\tanh$ or $g=\text{sigmoid}$. This implies that $\mathbf{Z}(t)$ contains an abstract representation of $\mu(t)$ [59].
@@
If the readout function $g$ is $\tanh$, then the reconstruction equation for $\mu(t)$ from $\mathbf{Z}(t)$ can be simplified using the approximation $\Phi(z)\approx\frac{1}{2}\tanh(z)+\frac{1}{2}$. Consequently, $\mu(t)$ can be expressed directly in terms of $\mathbf{Z}(t)$ without the need for the inverse CDF.
@@
The normal distribution CDF $\Phi(z)$ and the function $\frac{1}{2}\tanh(z)+\frac{1}{2}$ are known to be very similar, as both functions are sigmoid-like, centered at zero, and asymptotically approach 0 and 1 [14].
Page [60] proposed a simple approximation of $\Phi$ via $\tanh$. Eliminating higher order terms, their approximation is $\Phi(x)\approx\frac{1}{2}\tanh(\sqrt{\frac{2}{\pi}}x)+\frac{1}{2}$. We found the following approximation yielded a superior mean squared error:
@@
$\mu(t)\approx\frac{a_{\sigma}\,\sigma}{\sqrt{t}}(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\mathbf{C}^{\top}\mathbf{Z}(t)+(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\mathbf{b}$ (25)
@@
We found $a_{\sigma}=0.5886$ by computationally minimizing the mean squared error between $\Phi(a_{\sigma}z)$ and $\sigma(z)$ (see sigmoid_approx_gaussianCDF.m in supporting code). Upon computing $a_{\sigma}$ on successively larger optimization bounds $T=1,10,100,...$, we found that $a_{\sigma}$ converged to $0.5886$. We note that sigmoid approximations to the Gaussian distribution CDF have existed in the literature for some time [74]. ∎
@@
Corollary B.7 and B.8 are visualized in Figure S7, showing the close approximations to the Gaussian CDF.
@@
Recall equation (3) from the disentangled representation theorem:
$\mu(t)=(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}~\begin{pmatrix}\frac{\sigma}{\sqrt{t}}~\Phi^{-1}\big{(}g(\mathbf{Z}(t))\big{)}+\mathbf{b}\end{pmatrix}$
@@
$B^{\top}B$ is a symmetric for any matrix B. Recall that the inverse of a symmetric matrix is also symmetric. So $(\mathbf{C}^{\top}\mathbf{C})^{-1}$ is also symmetric. Therefore
@@
As the columns of $\mathbf{C}$ are high-dimensional randomly sampled vectors, their probability of being orthogonal vanishes as the dimensionality $N_{\textit{task}}$ increases. We can also state the condition in terms of the singular value decomposition (SVD) of $\mathbf{C}=\mathbf{U}\Sigma\mathbf{V}^{T}$ where $\mathbf{U}\in\mathbb{R}^{N_{\textit{task}}\times D}$, $\Sigma=\text{diag}(\sigma_{1},...,\sigma_{D})$, $\mathbf{V}\in\mathbb{R}^{D\times D}$, and $\sigma_{i}$ is the $i$th singular value of $\mathbf{C}$ and $\mathbf{U},\mathbf{V}$ are orthonormal. Then
@@
Therefore, uniform singular values in $\mathbf{C}$ is a sufficient condition to guarantee an orthogonal, disentangled representation of $\mu(t)$ in $\mathbf{Z}(t)$ [^4] This uniformity condition in the singular value decomposition is analogous to the outcome of the LM damping technique [45, 50], used for least squares inversion problems in various applications.. ∎

### v3

@@ Appendix B Theoretical Derivations
High-level summary of proof
We prove that competence at $N_{\text{task}}$ tasks guarantees linear decodability when $N_{\text{task}}\geq D$ for non-degenerate tasks (Theorem B.5), and orthogonality when $N_{\text{task}}\gg D$ and task boundaries are sampled randomly (Corollary B.10). To that end, we first show that optimal evidence aggregation in a multi-task classification framework enforces the multi-task classifier to encode a notion of distances from classification lines (Lemma B.3). Given a suitable set of distances from classification lines, we show that one can uniquely identify an optimal estimate of the latents given noisy data in closed form (Trilateration Theorem, Theorem B.4). In addition, if the readout function of the multi-task classifier is sigmoid-like, the optimal estimate will be approximately linearly decodable from the representation (Corrolary B.8,B.9). We then prove by contradiction that all the above results hold when an arbitrary injective observation map $f$ is applied to the input after noising (Theorem B.6). The theory generalizes to sub-optimal classifiers via least-mean squares approximation and the Moore-Penrose Pseudoinverse (Corollary B.7), and to different noise distributions (Section B.4). Finally, we discuss implications of the theorem for representation learning, manifold learning and the Platonic representation hypothesis, and future directions (Section B.6).
@@
  • $\hat{\mathbf{Y}}(t):=g(\mathbf{Z}(t))\in[0,1]^{N_{\textit{task}}}:$ Output vector of the multi-task classification model at time $t$, where each $\hat{Y}_{i}(t)$ is a Bernoulli random variable estimator, estimating the conditional probability $\Pr\{y_{i}(\mathbf{x}^{*})=1\}$ given the noisy observations (via latent variable $\mathbf{Z}(t)$ – see Equation 12).
@@ B.2 Trilateration via Multiple Decision Boundaries
: We derived an optimal estimator of $\mathbf{x}^{*}$ (denoted $\hat{\mathbf{X}}(t)$) based on noisy i.i.d. measurements $\mathbf{X}(1),\dots,\mathbf{X}(t)\sim\mathcal{N}(\mathbf{x}^{*},\sigma^{2}\mathbf{I}_{D})$ in Lemma B.1. In Lemma B.2 we derived the equation for Bernoulli variable estimator $\hat{Y}(t)$ to estimate a single classification output $y(\mathbf{x}^{*})$ based on the same noisy measurements via $\hat{\mathbf{X}}(t)$. Finally, we showed in Lemma B.3 that the uncertainty in $\hat{Y}(t)$ and the time $t$ is sufficient to determine the projection distance between the decision boundary and $\mu(t)=\text{mean}(\mathbf{X}(1),\dots,\mathbf{X}(t))$ via Equation 15.
Let $\hat{\mathbf{Y}}(t)$ denote the vector of classification estimates $\hat{Y}(t)$ from Equation 15. We now have the tools to prove our final result via trilateration. Much like distance information from cell towers can be used to trilaterate [^6] Trilateration differs from triangulation, and it is more frequently used in practice. Triangulation is when one has angle information w.r.t. the cell towers. Usually, this is not available – so one trilaterates their position Oguejiofor et al. (2013). This more closely matches our setting, where we just have distances information w.r.t. the decision boundaries and must determine the position. one’s position, we will leverage Lemma B.3 and use distances from decision boundaries $\{(\mathbf{c}_{i},b_{i})\}_{i\in[N_{\textit{task}}]}$ to constrain the positions.
@@
If $\mathbf{C}\in\mathbb{R}^{N_{task}\times D}$ is full-rank and $N_{task}\geq D$, then $\hat{\mathbf{Y}}(t)$, $t$, $\mathbf{b}$, and $\mathbf{C}$ are sufficient to reconstruct the exact value of $\mathbf{\mu}(t)$, the mean of $\mathbf{X}(1),\dots,\mathbf{X}(t)$, which is also the optimal estimator for $\mathbf{x}^{*}$.
@@
We may prove this claim by providing an algorithm to reconstruct $\mathbf{\mu}(t)=\text{mean}(\mathbf{X}(1),\dots,\mathbf{X}(t))$ from $\hat{\mathbf{Y}}(t),\mathbf{C}$, and $t$. Invoke Lemma B.3 to compute the signed projection distance between $\mathbf{\mu}(t)$ and each decision plane $(\mathbf{c}_{i},b_{i})$. Let $\mathbf{k}=[k_{1},\dots,k_{N_{task}}]^{\top}$ where each $k_{i}$ corresponds to decision boundary $\mathbf{c}_{i}$. Then the mean $\mathbf{\mu}(t)$ must satisfy
@@
Thus, for full rank $\mathbf{C}$ and $N_{task}\geq D$, we will have a uniquely determined $\mu(t)$ value. ∎
Sufficient statistics and optimal estimators:
“A statistic $\mu(t)$ is called sufficient for $\mathbf{x}^{*}$ if it contains all the information in ${\mathbf{X}(1),\dots,\mathbf{X}(t)}$ about $\mathbf{x}^{*}$.” (from Cover and Thomas’ Elements of Information Theory, 1999, Section 2.10, substituting variable names).
More formally, “A function $T({\mathbf{X}(1),\dots,\mathbf{X}(t)})$ is said to be a sufficient statistic relative to the family [of probability density functions indexed by $\mathbf{x}^{*}$] $f({\mathbf{X}(1),\dots,\mathbf{X}(t)}|\mathbf{x}^{*})$ if ${\mathbf{X}(1),\dots,\mathbf{X}(t)}$ is independent of $\mathbf{x}^{*}$ given $T({\mathbf{X}(1),\dots,\mathbf{X}(t)})$, i.e. $\mathbf{x}^{*}\to T\big{(}{\mathbf{X}(1),\dots,\mathbf{X}(t)}\big{)}\to\mathbf{X}(1),\dots,\mathbf{X}(t)$ forms a Markov chain. This is the same as the condition for equality in the data processing inequality,
$I(\mathbf{x}^{*};{\mathbf{X}(1),\dots,\mathbf{X}(t)})=I(\mathbf{x}^{*};\mu(t))$
for all distributions on $\mathbf{x}^{*}$. Hence sufficient statistics preserve mutual information and conversely.” (Cover and Thomas’ Elements of Information Theory, 1999, Section 2.10, substituting variable names)
$\mu(t)=\text{mean}(\mathbf{X}(1),\dots,\mathbf{X}(t))$ is a sufficient statistic for $\mathbf{x}^{*}$ given measurements $\mathbf{X}(i)\sim\mathbf{x}^{*}+\sigma\mathcal{N}(0,\mathbf{I}_{D})$: For Gaussian noise it’s a well known result that the sufficient statistic for the underlying mean given i.i.d. samples is the sample mean of the observations (Cover and Thomas, Elements of Information Theory, 1999, Section 2.10).
@@
Any system that optimally estimates classification probabilities $\hat{\mathbf{Y}}(t)$ based on noisy measurements $\{\mathbf{X}(1),\dots,\mathbf{X}(t)\}$ must implicitly encode a representation of $\mu(t)=\text{mean}(\mathbf{X}(1),\dots,\mathbf{X}(t))$ in its latent state $\mathbf{Z}(t)$ if decision boundary matrix $\mathbf{C}$ is full rank and $N_{task}\geq D$.
@@
We showed in Theorem B.4 (Trilateration Theorem) that if $\mathbf{C}\in\mathbb{R}^{N_{task}\times D}$ is full-rank and $N_{task}\geq D$, then $\hat{\mathbf{Y}}(t)$, $t$, $\mathbf{b}$, and $\mathbf{C}$ are sufficient to reconstruct the exact value of $\mu(t)$, the mean of $\mathbf{X}(1),\dots,\mathbf{X}(t)$. Rearranging Equation 16 and applying Equation 15,
$\mu(t)=(\mathbf{C^{\top}C})^{-1}\mathbf{C}^{\top}(\frac{\sigma}{\sqrt{t}}\Phi^{-1}(\hat{\mathbf{Y}}(t))+\mathbf{b})$
Replacing $\hat{\mathbf{Y}}(t)=g(\mathbf{Z}(t))$ from our problem setup reveals that $\mu(t)$ is a deterministic function of $\mathbf{Z}(t)$.
Therefore, optimal multi-task classifier latent state $\mathbf{Z}(t)$ contains a sufficient statistic $\mu(t)$ of $\mathbf{x}^{*}$, which implies that $\mathbf{Z}(t)$ must also contain all information about $\mathbf{x}^{*}$ given noisy measurements ${\mathbf{X}(1),\dots,\mathbf{X}(t)}$ if $\mathbf{C}\in\mathbb{R}^{N_{task}\times D}$ is full-rank and $N_{task}\geq D$. ∎
@@
$\tilde{\mu}=(\mathbf{C^{\top}C})^{-1}\mathbf{C}^{\top}(\mathbf{k}+\mathbf{b})$ (17)
Conveniently, if the estimation errors in sub-optimal $\hat{\mathbf{Y}}$ have a mean of zero, additional decision boundaries in $\mathbf{C}$ (e.g., beyond the minimum $D$ linearly independent boundaries) result in improved estimation of $\mathbf{x}^{*}$ by the central limit theorem, thus generalizing our results to sub-optimal estimators (see Corollary B.7).
@@ B.3 Optimal Representation Theorem (General Case)
If $\mathbf{C}\in{N_{\textit{task}}\times D}$ is a full-rank matrix representing the decision boundary normal vectors in $\mathbb{R}^{D}$ and $N_{\textit{task}}\geq D$, then any optimal estimator of $\mathbf{y}(\mathbf{x}^{*})$ must encode an optimal estimator $\mu(t)$ of the latent variable $\mathbf{x}^{*}$ in its latent state $\mathbf{Z}(t)$. Furthermore, $\mu(t)$ is a sufficient statistic of $\mathbf{x}^{*}$, ensuring that all the information about $\mathbf{x}^{*}$ contained in $\{\mathbf{X}(t)\}$ is also contained in $\mathbf{Z}(t)$. Consequently, $\mu(t)$ – the optimal estimate of $\mathbf{x}^{*}$ based on $f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))$ – can be written as a deterministic function (Equation 18) of latent state $\mathbf{Z}(t)$.
@@
[…]
@@
where $g(\mathbf{Z}(t))$ represents the transformation from the latent state to the classification probabilities.
Since $\mu(t)$ is the optimal estimator (and sufficient statistic) for $\mathbf{x}^{*}$ given measurements $\{\mathbf{X}(1),\dots,\mathbf{X}(t)\}$, it contains all information about $\mathbf{x}^{*}$ contained in the measurements (Cover & Thomas (1991)). In other words, $\mu(t)$ is a deterministic function of $\mathbf{Z}(t)$, implying that $\mathbf{Z}(t)$ will contain all information about $\mathbf{x}^{*}$ contained in the measurements $\{\mathbf{X}(t)\}$. ∎
Corollary B.7 (Recovery of $\mu(t)$ for Sub-Optimal Classifiers).
Let $\hat{\mathbf{Y}}(t)\in[0,1]^{N_{\text{task}}}$ represent the output of a sub-optimal classifier with zero-mean independent errors, i.e., $\hat{Y}_{i}(t)=\Pr\{y_{i}(\mathbf{x^{*}})=1\}+\epsilon_{i}$, where $\mathbb{E}[\epsilon_{i}]=0$ and $\text{Var}[\epsilon_{i}]=\sigma_{\epsilon}^{2}$ for all $i\in\{1,\dots,N_{\text{task}}\}$.
If $\mathbf{C}\in\mathbb{R}^{N_{\text{task}}\times D}$ is a full-rank and well-conditioned matrix of decision boundary normal vectors with $N_{\text{task}}\geq D$, the estimated mean $\tilde{\mu}(t)$ of $\mathbf{x^{*}}$ can be recovered using the Moore-Penrose pseudoinverse:
$\tilde{\mu}(t)=(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}(\mathbf{k}+\mathbf{b}),$
where $\mathbf{k}=\frac{\sigma}{\sqrt{t}}\Phi^{-1}(\hat{\mathbf{Y}}(t))$ and $\Phi^{-1}$ is the inverse CDF of the standard normal distribution.
For sub-optimal classifiers, as the number of tasks $N_{\text{task}}$ increases:
  • The redundancy in $\mathbf{C}$ reduces sensitivity to classification errors.
  • Under the assumption of independent, zero-mean errors in $\hat{\mathbf{Y}}(t)$, the residual error in $\tilde{\mu}(t)$ is expected to decrease at a rate of approximately $\mathcal{O}(1/\sqrt{N_{\text{task}}})$, driven by the averaging effect of least-squares estimation.
Motivated by the similarity between $\Phi(z)$ and sigmoid-like activation functions $g(z)$, we show that the two can be approximately canceled in Equation 18, implying that $\mu(t)$ can be reconstructed with high accuracy with a linear-affine transformation (e.g., linear decoding) when $g=\tanh$ or $g=\text{sigmoid}$. This implies that $\mathbf{Z}(t)$ contains an abstract representation of $\mu(t)$ (Ostojic & Fusi, 2024).
@@
If the readout function $g$ is $\tanh$, then the reconstruction equation for $\mu(t)$ from $\mathbf{Z}(t)$ can be simplified using the approximation $\Phi(z)\approx\frac{1}{2}\tanh(\frac{\pi}{2\sqrt{3}}z)+\frac{1}{2}$. Consequently, $\mu(t)$ can be expressed directly in terms of $\mathbf{Z}(t)$ without the need for the inverse CDF.
@@
The normal distribution CDF $\Phi(z)$ and the function $\frac{1}{2}\tanh(z)+\frac{1}{2}$ are known to be very similar, as both functions are sigmoid-like, centered at zero, and asymptotically approach 0 and 1 (Choudhury, 2014).
Page (1977) proposed a simple approximation of $\Phi$ via $\tanh$. Eliminating higher order terms, their approximation is $\Phi(x)\approx\frac{1}{2}\tanh(\sqrt{\frac{2}{\pi}}x)+\frac{1}{2}$. We found the following approximation yielded a superior mean squared error:
@@
$\mu(t)\approx\frac{a_{\sigma}\,\sigma}{\sqrt{t}}(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\mathbf{Z}(t)+(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\mathbf{b}$ (22)
@@
We found $a_{\sigma}=0.5886$ by computationally minimizing the mean squared error between $\Phi(a_{\sigma}z)$ and $\sigma(z)$ (see sigmoid_approx_gaussianCDF.m in supporting code). Upon computing $a_{\sigma}$ on successively larger optimization bounds $T=1,10,100,...$, we found that $a_{\sigma}$ converged to $0.5886$. We note that sigmoid approximations to the Gaussian distribution CDF have existed in the literature for some time (Waissi & Rossin, 1996). ∎
@@
Corollary B.8 and B.9 are visualized in Figure S9, showing the close approximations to the Gaussian CDF.
@@
Recall Equation 3 from the disentangled representation theorem:
[…]
@@
$B^{\top}B$ is a symmetric matrix for any matrix B. Recall that the inverse of a symmetric matrix is also symmetric. So $(\mathbf{C}^{\top}\mathbf{C})^{-1}$ is also symmetric. Therefore
@@
As the columns of $\mathbf{C}$ are high-dimensional randomly sampled vectors, their probability of being non-orthogonal vanishes as the dimensionality $N_{\textit{task}}$ increases. We can also state the condition in terms of the singular value decomposition (SVD) of $\mathbf{C}=\mathbf{U}\Sigma\mathbf{V}^{T}$ where $\mathbf{U}\in\mathbb{R}^{N_{\textit{task}}\times D}$, $\Sigma=\text{diag}(\sigma_{1},...,\sigma_{D})$, $\mathbf{V}\in\mathbb{R}^{D\times D}$, and $\sigma_{i}$ is the $i$th singular value of $\mathbf{C}$ and $\mathbf{U},\mathbf{V}$ are orthonormal. Then
@@
Therefore, uniform singular values in $\mathbf{C}$ is a sufficient condition to guarantee an orthogonal, disentangled representation of $\mu(t)$ in $\mathbf{Z}(t)$ [^7] This uniformity condition in the singular value decomposition is analogous to the outcome of the LM damping technique (Levenberg, 1944; Marquardt, 1963), used for least squares inversion problems in various applications.. ∎
As a sidenote, we would like to point out that the setting here relates to the Marchenko-Pastur law while noting important caveats: while the law typically applies to matrices with i.i.d. entries $N(0,\sigma)$, our $\mathbf{C}$ matrix consists of random row vectors in $\mathbb{R}^{D}$ of unit norm. This structure, while not strictly meeting the law’s conditions, still supports our conclusions about orthogonalization.